Amount of an investment. The amount of an investment of P dollars for t years at simple interest rate r is given by A=P + Prt.
a) rewrite this formula by factoring out the greatest common factor on the right hand side.
b) Find A if $8300 is invested for 3 years at a simple interest rate of 15%.

Answer by rfer(16322)

A=P(1+r)t
A=8300(1+.15)3
A=8300*3.45
A=$28,635